OpenStudy (anonymous):

The shortest distance from the curve xy=4 to the origin is...

OpenStudy (amistre64):

the point at which the slope of the line is perpendicular to a line thru the origin

OpenStudy (anonymous):

Is this calculus?

OpenStudy (anonymous):

sqrt(8) The distance between some point of the curve and the origin can be calculated in such a way \[\sqrt{x ^{2}+(f(x))^2}\] We need to find the minimum of that function. Let's look at it's derivative: \[(\sqrt{x^2+(f(x))^2})'=\dfrac{x^2-16/x^2}{\sqrt(x^4+16)}\] The derivative equals zero at one point x=2/ It is the minimum as the sign changes from - to +. So the point we are tryinf to find is x=2 y=2. So the distanse is sqrt(8).

OpenStudy (anonymous):

f(x)=y=4/x

OpenStudy (anonymous):

ya differentiaion apparenly

OpenStudy (anonymous):

If so, \[x^2+y^2 ---> to \space minimize\] constraint \[xy=4\] \[\nabla f=\nabla g \lambda\] \[\nabla f=<2x,2y>\] \[\nabla g=<y,x>\] \[2x=y\lambda\] \[2y=x\lambda\] \[xy=4\] x=2 y=2 Sqrt[2^2 +2^2]=Sqrt[8]